### v1

# A Necessary Condition for the Submergence of Proto-Neutron Star Magnetic Fields by Supernova Fallback

Abstract
Central compact objects (CCOs) are a subclass of neutron stars with a dipole magnetic field strength considerably weaker than those of radio pulsars and magnetars. One possible explanation for such weak magnetic fields in the CCOs is the hidden magnetic field scenario, in which supernova fallback submerges the magnetosphere of a proto-neutron star beneath a newly formed crust. However, the fallback mass and timescale required for this submergence process remain uncertain. We perform one-dimensional general relativistic magnetohydrodynamic simulations of the supernova fallback onto a magnetized proto-neutron star, while considering neutrino cooling. In our simulations, the infalling material compresses the magnetic field and drives a strong shock. The shock initially expands outward, but eventually stalls and recedes as neutrino cooling becomes significant. After the shock stalls, the gas density above the magnetosphere increases rapidly, potentially leading to the formation of a new crust. To understand the shock dynamics, we develop semi-analytic models that describe the resulting magnetospheric and shock radii when the shock stalls. By comparing the fallback time scale with the shock stalling time scale, corresponding to the waiting time for the new crust formation, we derive a necessary condition for the submergence of the PNS magnetic field. Our results will provide guidance for investigating the diversity of young isolated neutron stars through multidimensional simulations.
Neutron stars (1108) — General relativity (641) — Accretion (14) — High energy astrophysics (739)
{CJK*}
UTF8ipxm
1 Introduction
Based on multi-wavelength observations, young isolated neutron stars (NSs) in our galaxy with ages of $\lesssim 1-10~{}{\rm kyr}$ are classified as magnetars, radio pulsars, and central compact objects (CCOs). The central engines powering these objects are believed to be the magnetic energy, rotation energy, and thermal energy of the NSs, respectively (e.g., see reviews by Enoto et al., 2019; Borghese & Esposito, 2023). Their characteristic values of the dipole magnetic field strength at the NS surface are considerably different: $>10^{14}~{}{\rm G}$ for magnetars, $10^{12-13}~{}{\rm G}$ for radio pulsars, and $<10^{11}~{}{\rm G}$ for CCOs. However, the origin of such diversity remains unclear.
What determines the magnetic field strength of NSs is an open question (e.g., see reviews by Igoshev et al., 2021). When a proto-NS (PNS) forms in a core-collapse supernova, the magnetic field of the progenitor core is expected to undergo amplification through various mechanisms: magnetic flux conservation (Woltjer, 1964), turbulent dynamo action (Duncan & Thompson, 1992; Thompson & Duncan, 1993; Masada et al., 2022), magnetorotational instability inside the PNS (Obergaulinger et al., 2009; Akiyama et al., 2003; Masada et al., 2015), and/or steady accretion shock instability of the post-bounce core-collapse supernova environment (Blondin et al., 2003). On the other hand, it is thought that the NS magnetic field decays over a relatively long time scale, $\gg 10~{}{\rm kyr}$, due to Ohmic diffusion, ambipolr diffusion, and Hall drift (e.g., Goldreich & Reisenegger, 1992; Takahashi et al., 2011; Skiathas & Gourgouliatos, 2024).
Supernova fallback has also been proposed as a contributing factor in the magnetic field evolution, particularly in the formation of the CCOs (see e.g., Muslimov & Page, 1995; Young & Chanmugam, 1995; Geppert et al., 1999). The fallback phase starts several seconds after the core bounce at the time when the neutrino-driven wind subsides (see e.g., Janka et al., 2022). If the fallback accretion rate is sufficiently high, the PNS’s magnetic field will be submerged beneath a crust newly formed by the fallback material. This case corresponds to the CCO formation (Torres-Forné et al., 2016; Shigeyama & Kashiyama, 2018). This scenario is referred to as the hidden magnetic field scenario. The hidden magnetic field is also believed to be the energy source of giant flares observed in soft gamma-ray repeaters (Thompson & Duncan, 2001).
The fallback accretion rate remains highly uncertain. Based on numerical simulations of the supernova explosion (e.g., Zhang et al., 2008; Dexter & Kasen, 2013; Moriya et al., 2018), the fallback accretion rate is approximated to be (Metzger et al., 2018; Barrère et al., 2022)
$\dot{M}_{\rm fb}=\frac{2}{3}\frac{M_{\rm fb}}{t_{\rm fb}}\left(\frac{1}{1+t/t_{\rm fb}}\right)^{5/3},$ (1)
where $t$ stands for time, and $M_{\rm fb}$ and $t_{\rm fb}$ are the typical fallback mass and fallback time scale, respectively. The dependence of $\dot{M}_{\rm fb}\propto t^{-5/3}$ for $t\gg t_{\rm fb}$ is expected for an accretion of marginally gravitationally bound matter (Michel, 1988; Chevalier, 1989). The typical fallback mass is estimated to be $M_{\rm fb}\sim 10^{-(4-1)}~{}{\rm M_{\odot}}$, depending on the progenitor structure (see e.g., Ugliano et al., 2012; Ertl et al., 2016a, b). The fallback time $t_{\rm fb}$ typically ranges from $1~{}{\rm s}$ to $10^{3}~{}{\rm s}$ (Janka et al., 2022), depending on the mean density of the progenitor star layer from which the fallback mass $M_{\rm fb}$ originates (Metzger et al., 2018). Therefore, $\dot{M}_{\rm fb}$ spans a wide range of $(10^{-7}-10^{-1})~{}{\rm M_{\odot}~{}s^{-1}}$.
Figure 1: A schematic view of the supernova fallback. The white circle in the left figure denotes the PNS with a radius of $r_{\rm PNS}$. gray arrows stand for the velocity vector of the free-falling gas, and black wavy lines with arrows indicate the neutrinos propagating outward. We denote the magnetospheric and shock radii as $r_{\rm M}$ and $r_{\rm sh}$, respectively, in terms of the spherical radius $r$. The overall structure of the supernova fallback can be divided into free-fall region ($r>r_{\rm sh}$), post-shock region ($r_{\rm M}<r<r_{\rm sh}$), and the PNS magnetosphere ($r_{\rm PNS}<r<r_{\rm M}$). The shadow in the post-shock region represents the gas density and temperature. The right two panels are enlarged views around $r=r_{\rm M}$. A new crust forms above the PNS magnetosphere when neutrino cooling is efficient (panel a), whereas no crust forms when neutrino cooling is inefficient (panel b).
As illustrated in the left panel of Figure 1, we consider that the fallback accretion onto a magnetized PNS consists of three regions (Torres-Forné et al., 2016): the free-fall region ($r>r_{\rm sh}$), the post-shock region ($r_{\rm M}<r<r_{\rm sh}$), and the PNS magnetosphere ($r_{\rm PNS}<r<r_{\rm M}$). Here, we denote the magnetospheric and shock radii by $r_{\rm M}$ and $r_{\rm sh}$, respectively, in terms of the spherical radius $r$. We write the PNS radius as $r_{\rm PNS}$. For $r>r_{\rm sh}$, the gas accretes toward the PNS at a free-fall velocity. For $r_{\rm M}<r<r_{\rm sh}$, a strong shock forms and the fallback material slowly settles to the surface of $r=r_{\rm M}$. For $r_{\rm PNS}<r<r_{\rm M}$, the PNS’s magnetic field is confined by the post-shock flows. When $\dot{M}_{\rm fb}$ is very high, neutrino emission occurs at $r\sim r_{\rm M}$. Such neutrino emission releases the gravitational energy liberated at $r\sim r_{\rm M}$. The two panels on the right are enlarged views around $r=r_{\rm M}$. These panels describe the cases where neutrino cooling is effective (panel a) and ineffective (panel b). If the neutrino cooling is effective enough for the new crust to form, then the PNS magnetosphere is submerged. Additionally, if the PNS rotation is significant, a bipolar jet powered by the magnetic field (Parfrey & Tchekhovskoy, 2017) or neutrinos (Qian & Woosley, 1996) will emerge around the rotation axis.
Neutrino emissions affect the dynamics of the supernova fallback. Chevalier (1989) found steady solutions in which the gravitational energy released at the PNS surface is lost through neutrino cooling in the post-shock region. The steady solution is later confirmed by one-dimensional general relativistic Boltzmann neutrino-hydrodynamic simulations by Akaho et al. (2024). Since these studies considered the non-magnetzed PNS, it is still not understood how the neutrino emission affects the dynamics of the supernova fallback onto a magnetized PNS.
In three dimensions, convection can influence the submergence of the PNS magnetic fields through convective diffusion. Bernal et al. (2013) presents a set of three-dimensional magnetohydrodynamic (MHD) simulations examining the interaction between fallback matter and an isolated magnetic arcade, suggesting that the post-shock region is marginally convectively stable (see also Bernal et al., 2010). Therefore, one- and two-dimensional models that neglect the effects of convection remain reasonable. Since this pioneering three-dimensional study explored only a limited parameter space, comprehensive surveys using one- or two-dimensional models are important for identifying the conditions under which the submergence of the magnetic field occurs.
In this study, we perform one-dimensional general relativistic MHD (GRMHD) simulations of the supernova fallback onto a magnetized PNS over wide parameter ranges. Our simulations take neutrino cooling into account. We study how the neutrino cooling affects the dynamics of fallback material for various PNS magnetic field strengths. Using numerical models, we derive a necessary condition for the submergence of the PNS magnetic field.
This paper is organized as follows. In Section 2, we briefly review the steady solution without magnetic fields in Chevalier (1989), which provides a key concept of the fallback process. We will present the numerical methods and models in Section 3 and show the numerical results in Section 4. Section 5 is devoted to the discussion of the diversity of young isolated NSs (Section 5.1) and neutrino scattering and absorption (Section 5.2). Finally, we give our conclusion in Section 6. Hereafter, the speed of light $c$ and the gravitational constant $G$ are normalized to $1$ unless otherwise specified. We represent the PNS mass as $M_{\rm PNS}$ and the gravitational radius as $r_{\rm g}=M_{\rm PNS}$.
2 Overview of Chevalier 1989
We briefly review the steady solution of the supernova fallback onto a non-magnetized PNS (Chevalier, 1989), which provides a guideline of our analysis. Since neutrino cooling is effective only in the vicinity of the PNS surface, the post-shock flow will be adiabatic over most of the accretion flow. We consider situations in which radiation pressure dominates gas pressure because of high accretion rates. To take into account the effect, we take an effective specific heat ratio of $\Gamma=4/3$ for the equation of state. Additionally, since local thermodynamic equilibrium will be achieved in the flows, we regard the radiation and gas temperatures as equal. Thus, we calculate the gas temperature as $T=(3p/a_{\rm rad})^{1/4}$, where $a_{\rm rad}$ is the radiation constant.
From the Rankine-Hugoniot relations, the gas density $\rho$, pressure $p$, and velocity $v$ in the post-shock region for a standing shock at the radius of $r_{\rm sh}$ are described as functions of radius $r$:
$\rho(r)$ $=$ $7\rho_{\rm pre}\left(\frac{r}{r_{\rm sh}}\right)^{-3},$ (2)
$p(r)$ $=$ $\frac{6}{7}\rho_{\rm pre}v_{\rm ff}^{2}(r_{\rm sh})\left(\frac{r}{r_{\rm sh}}\right)^{-4},$ (3)
$v(r)$ $=$ $\frac{v_{\rm ff}(r_{\rm sh})}{7}\left(\frac{r}{r_{\rm sh}}\right),$ (4)
where $\rho_{\rm pre}=\dot{M}_{\rm fb}/(4\pi r_{\rm sh}^{2}v_{\rm ff}(r_{\rm sh}))$ is the pre-shock gas density at $r=r_{\rm sh}$, and $v_{\rm ff}$ is the free-fall velocity from infinity.
The steady-state solution is obtained by assuming that all gravitational energy released by accretion is lost through the neutrino emission:
$\frac{M_{\rm PNS}\dot{M}_{\rm in}}{r_{\rm PNS}}=4\pi r_{\rm PNS}^{2}H\dot{q},$ (5)
where $\dot{q}$ is the cooling rate due to the neutrino emission. Chevalier (1989) adopted $H=r_{\rm PNS}^{2}p/({M_{\rm PNS}\rho})=(12/49)r_{\rm PNS}$ for the fiducial width of the layer in which neutrino cooling works effectively. From Equations (2) - (5), we get $r_{\rm sh}$ as a function of $\dot{M}_{\rm fb}$.
Figure 2: Left panel: cooling rate due to the neutrino emission against the gas temperature. We employ $\rho=10^{6},~{}10^{8},~{}10^{10}$, and $10^{12}~{}{\rm cm^{3}~{}s^{-1}}$. Right panel: steady solutions of the shock radius as a function of fallback accretion rate. Here, we adopt $r_{\rm PNS}=10~{}{\rm km}$ and $M_{\rm PNS}=1.4M_{\odot}$ (see Equations 8 and 9). The vertical dashed line denotes the mass accretion rate when $r_{\rm sh,pair}=r_{\rm sh,URCA}$ ($\dot{M}_{\rm fb,a}$).
The left panel of Figure 2 shows the cooling rate as a function of the gas temperature. We consider two processes relevant to the fallback accretion (Itoh et al., 1989; Qian & Woosley, 1996):
$\dot{q}_{\rm pair}$ $=$ $5\times 10^{33}[{\rm erg~{}cm^{-3}~{}s^{-1}}]\left(\frac{T}{10^{11}~{}{\rm K}}\right)^{9}.$ (6)
$\dot{q}_{\rm URCA}$ $=$ $9\times 10^{29}[{\rm erg~{}cm^{-3}~{}s^{-1}}]\left(\frac{\rho}{10^{6}~{}{\rm g~{}cm^{-3}}}\right)\left(\frac{T}{10^{11}~{}{\rm K}}\right)^{6},$ (7)
The pair neutrino process is $e^{-}+e^{+}\rightarrow\nu_{\rm e}+\bar{\nu}_{\rm e}$, and the URCA processes are $p+e^{-}\rightarrow n+\nu_{\rm e}$ and $n+e^{+}\rightarrow p+\bar{\nu}_{\rm e}$. For $\rho=10^{8}~{}{\rm g~{}cm^{-3}}$, $\dot{q}_{\rm pair}$ is dominant over $\dot{q}_{\rm URCA}$ for $T\gtrsim 3\times 10^{10}~{}{\rm K}$.
The right panel of Figure 2 displays the analytical solutions of the shock radius $r_{\rm sh,pair}$ (gold line) and $r_{\rm sh,URCA}$ (black line) against $\dot{M}_{\rm fb}$:
$r_{\rm sh,pair}$ $=$ $3.4\times 10^{2}~{}[{\rm km}]\left(\frac{r_{\rm PNS}}{10~{}{\rm km}}\right)^{40/27}\left(\frac{M_{\rm PNS}}{1.4~{}{\rm M}_{\odot}}\right)^{-1/27}\left(\frac{\dot{M}_{\rm fb}}{10^{-5}~{}{\rm M_{\odot}~{}s^{-1}}}\right)^{-10/27},$ (8)
$r_{\rm sh,URCA}$ $=$ $4.3\times 10^{2}~{}[{\rm km}]\left(\frac{r_{\rm PNS}}{10~{}{\rm km}}\right)^{4/3}\left(\frac{M_{\rm PNS}}{1.4~{}{\rm M}_{\odot}}\right)^{1/5}\left(\frac{\dot{M}_{\rm fb}}{10^{-5}~{}{\rm M_{\odot}~{}s^{-1}}}\right)^{-2/5},$ (9)
for $\dot{q}=\dot{q}_{\rm pair}$ and $\dot{q}=\dot{q}_{\rm URCA}$, respectively. Because both $\dot{q}_{\rm pair}$ and $\dot{q}_{\rm URCA}$ increase with $\dot{M}_{\rm fb}$, both $r_{\rm sh,pair}$ and $r_{\rm sh,URCA}$ decrease with $\dot{M}_{\rm fb}$. The vertical dashed line indicates the accretion rate for $r_{\rm sh,pair}=r_{\rm sh,URCA}$ ($\dot{M}_{\rm fb,a}\approx 4\times 10^{-2}~{}{\rm M_{\odot}~{}s^{-1}}$). When $\dot{M}_{\rm fb}<\dot{M}_{\rm fb,a}$, we get the relation of $r_{\rm sh,pair}<r_{\rm sh,URCA}$. This relation indicates $\dot{q}_{\rm pair}>\dot{q}_{\rm URCA}$ for $\dot{M}_{\rm fb}<\dot{M}_{\rm fb,a}$. On the other hand, $\dot{q}_{\rm pair}<\dot{q}_{\rm URCA}$ holds for $\dot{M}_{\rm fb}>\dot{M}_{\rm fb,a}$. The shock radii are smaller than the PNS radius for $\dot{M}_{\rm fb}\gtrsim 10^{-1}~{}{\rm M_{\odot}~{}s^{-1}}$, indicating that the free-fall region reaches the PNS surface. In the present study, we investigate the supernova fallback with a mass accretion rate of $\dot{M}_{\rm fb}=10^{-(5-2)}~{}{\rm M_{\odot}~{}s^{-1}}$ (see Section 3.2 for details).
3 Numerical Method
We numerically solve the GRMHD equations while considering neutrino cooling in Schwarzschild polar coordinates $(t,r,\theta,\phi)$. These simulations are one-dimensional in spherical symmetry. We use the numerical code UWABAMI (e.g., Takahashi & Ohsuga, 2017). We represent space-time and space components as Greek and Latin suffixes, respectively.
3.1 Basic equations
The equations for the time-evolution of the GRMHD are as follows:
$\nabla_{\mu}\left(\rho u^{\mu}\right)$ $=$ $0,$ (10)
$\nabla_{\mu}\left(T^{\mu\nu}\right)$ $=$ $\sqrt{-g}Q^{\nu},$ (11)
$\partial_{t}\left(\sqrt{-g}B^{i}\right)$ $=$ $-\partial_{j}\left\{\sqrt{-g}\left(b^{i}u^{j}-b^{j}u^{i}\right)\right\},$ (12)
where $u^{\mu}$ is the four-velocity of the gas, $B^{i}$ is the magnetic field vector in the laboratory frame, $b^{\mu}$ is the magnetic four-vector in the fluid frame, and $g={\rm det}(g_{\mu\nu})$ is the determinant of the metric. The energy-momentum tensor of the ideal MHD is given by
${T}^{\mu\nu}$ $=$ $\left(\rho+e+p+\frac{b^{2}}{8\pi}+p_{\rm mag}\right)u^{\mu}u^{\nu}+\left(p+p_{\rm mag}\right)g^{\mu\nu}-\frac{b^{\mu}b^{\nu}}{4\pi},$ (13)
where $e=p/(\Gamma-1)$, $b^{2}=b^{\mu}b_{\mu}$, and $p_{\rm mag}=b^{2}/8\pi$ is the magnetic pressure in the fluid frame. We express the source term accounting for the neutrino emission as
$Q^{\mu}$ $=$ $-\left(\dot{q}_{\rm pair}+\dot{q}_{\rm URCA}\right)u^{\mu}.$ (14)
The numerical methods used to solve the above equations are as follows. We utilize a harmonic mean proposed by van Leer (1977) for the reconstruction and adopt the Harten–Lax–van Leer (HLL; Harten et al., 1983) scheme to evaluate the numerical flux. We perform the time integration of the advection term using an explicit second-order Runge-Kutta scheme. On the other hand, we implicitly solve the source term responsible for the neutrino cooling $Q^{\mu}$ with the Newton-Raphson method. Since the gas velocity in the region where neutrino cooling is pronounced is much smaller than the speed of light, we take $Q^{i}=0$. This approach enables us to solve the source term using the Newton-Raphson method for a single variable, $p$, alone, rather than for four variables, $p$ and $u^{i}$, thereby reducing computational cost. Although we do not solve the energy and momentum transport due to neutrino scattering and absorption, these processes do not affect our conclusions within the density range treated in the present study (see Section 5.2).
Table 1: Parameters for numerical models
Parameters $B_{\rm PNS}$ $\rho_{\rm out}$ $\dot{M}_{\rm fb}$ $t_{\rm shock}$
Unit $[\rm G]$ $[\rm g~{}cm^{-3}]$ $[\rm M_{\odot}~{}s^{-1}]$ $[\rm s]$
BnMm5R10 00 $10^{2}$ $10^{-5}$ $3.48\times 10^{-2}$
BnMm4R10 00 $10^{3}$ $10^{-4}$ $3.49\times 10^{-2}$
BnMm3R10 00 $10^{4}$ $10^{-3}$ $3.48\times 10^{-2}$
BnMm2R10 00 $10^{5}$ $10^{-2}$ $3.48\times 10^{-2}$
B13Mm5R10 $10^{13}$ $10^{2}$ $10^{-5}$ $3.49\times 10^{-2}$
B13Mm4R10 $10^{13}$ $10^{3}$ $10^{-4}$ $3.49\times 10^{-2}$
B13Mm3R10 $10^{13}$ $10^{4}$ $10^{-3}$ $3.49\times 10^{-2}$
B13Mm2R10 $10^{13}$ $10^{5}$ $10^{-2}$ $3.48\times 10^{-2}$
B14Mm5R10 $10^{14}$ $10^{2}$ $10^{-5}$ $3.54\times 10^{-2}$
B14Mm4R10 $10^{14}$ $10^{3}$ $10^{-4}$ $3.51\times 10^{-2}$
B14Mm3R10 $10^{14}$ $10^{4}$ $10^{-3}$ $3.50\times 10^{-2}$
B14Mm2R10 $10^{14}$ $10^{5}$ $10^{-2}$ $3.49\times 10^{-2}$
B15Mm4R10 $10^{15}$ $10^{3}$ $10^{-4}$ $3.61\times 10^{-2}$
B15Mm3R10 $10^{15}$ $10^{4}$ $10^{-3}$ $3.54\times 10^{-2}$
B15Mm2R10 $10^{15}$ $10^{5}$ $10^{-2}$ $3.51\times 10^{-2}$
BnMm4R20 00 $10^{3}$ $10^{-4}$ $3.49\times 10^{-2}$
BnMm3R20 00 $10^{4}$ $10^{-3}$ $3.49\times 10^{-2}$
BnMm2R20 00 $10^{5}$ $10^{-2}$ $3.48\times 10^{-2}$
B13Mm4R20 $10^{13}$ $10^{3}$ $10^{-4}$ $3.50\times 10^{-2}$
B13Mm3R20 $10^{13}$ $10^{4}$ $10^{-3}$ $3.50\times 10^{-2}$
B13Mm2R20 $10^{13}$ $10^{5}$ $10^{-2}$ $3.48\times 10^{-2}$
B14Mm4R20 $10^{14}$ $10^{3}$ $10^{-4}$ $3.59\times 10^{-2}$
B14Mm3R20 $10^{14}$ $10^{4}$ $10^{-3}$ $3.53\times 10^{-2}$
B14Mm2R20 $10^{14}$ $10^{5}$ $10^{-2}$ $3.50\times 10^{-2}$
B15Mm3R20 $10^{15}$ $10^{4}$ $10^{-3}$ $3.72\times 10^{-2}$
B15Mm2R20 $10^{15}$ $10^{5}$ $10^{-2}$ $3.59\times 10^{-2}$
Note. — The model names are shown in the first column: “BXX” means the magnetic field strength at the PNS surface of $10^{{\rm XX}}~{}{\rm G}$ (“Bn” is the case of $B_{\rm PNS}=0$), “MmX” denotes the fallback mass accretion rate of $10^{-{\rm X}}~{}{\rm M_{\odot}~{}s^{-1}}$, and “RXX” represents the PNS radius of ${\rm XX}~{}{\rm km}$. We present the magnetic field strength at the PNS surface ($B_{\rm PNS}$), the gas density at $r=r_{\rm out}$ ($\rho_{\rm out}$), the mass accretion rate at $r=r_{\rm out}$ ($\dot{M}_{\rm fb}=-4\pi r_{\rm out}^{2}\rho_{\rm out}u^{r}$), and the time when the accretion shock forms ($t_{\rm shock}$).
3.2 Numerical setup
We set the radius of the outer boundary to $r_{\rm out}=10^{3}~{}{\rm km}$. The computational domain consists of $[r_{\rm PNS},r_{\rm out}]$. The number of numerical grid points is $N_{r}=16384$. The radius of the grid exponentially increases with $e^{x_{a}}$, where $x_{a}=\ln r_{\rm PNS}+a\times dX$, $a$ is the radial grid index, and $dX=(\ln r_{\rm out}-\ln r_{\rm PNS})/N_{r}$. The minimum grid width is $\Delta r\approx 280~{}{\rm cm}$, which is much smaller than the pressure scale height near the PNS surface $\sim 10^{5}{\rm cm}$ for $\rho\sim 10^{10}~{}{\rm g~{}cm^{3}}$ and $T\sim 10^{11}~{}{\rm K}$.
Table 1 presents the magnetic field strength at the PNS surface ($B_{\rm PNS}$), the gas density at $r=r_{\rm out}$ ($\rho_{\rm out}$), the mass accretion rate at $r=r_{\rm out}$ ($\dot{M}_{\rm fb}=-4\pi r_{\rm out}^{2}\rho u^{r}$), and the time at which the accretion shock forms ($t_{\rm shock}$). We will provide an explanation of $t_{\rm shock}$ in Section 4.1.
[…]
We consider a magnetized and non-rotating PNS. The PNS initially has a magnetic field given by $B^{(\theta)}=B_{\rm dip}=B_{\rm PNS}(r/r_{\rm PNS})^{-3}$ and $B^{r}=B^{\phi}=0$. The parentheses in the indices of physical quantities denote the values in the static observer frame. This configuration models the fallback material accreting perpendicularly to the dipole magnetic field. We will discuss how the PNS rotation affects the submergence process in Section 5.1.
We initially set the atmosphere in an MHD equilibrium to maintain the radial profile of $B^{(\theta)}$. From the relation of $dp_{\rm mag}/dr=-M_{\rm PNS}\rho/r^{2}$, we can describe the initial gas density as $\rho=(3/4\pi)B_{\rm PNS}^{2}(r_{\rm PNS}/r_{\rm g})(r/r_{\rm PNS})^{-5}$. We set the initial pressure ratio to $p/p_{\rm mag}=0.01$. We also investigate the case of $B_{\rm PNS}=0$ for comparison. In this case, we take $\rho=10^{-4}\rho_{\rm out}(r/r_{\rm PNS})^{-1/(\Gamma-1)}$ and $p=10^{-4}\rho_{\rm out}[(\Gamma-1)/\Gamma](r_{\rm g}/r_{\rm PNS})(r/r_{\rm PNS})^{-\Gamma/(\Gamma-1)}$, which satisfy $dp/dr=-M_{\rm PNS}\rho/r^{2}$. In all models, we initially set $u^{i}=0$.
In the ghost cells inside $r_{\rm PNS}$, we fix $B^{i}$ to the initial value and set $\rho$ and $p$ according to the zero-gradient boundary condition. We adopt the mirror boundary condition for $u^{r}$. The electrical conductivity inside the PNS is sufficiently high that the electric field can be considered negligible (Igoshev et al., 2021). Therefore, we set the numerical flux in the radial component of Equation (12) to 0 at $r=r_{\rm PNS}$. From the outer boundary at $r=r_{\rm out}$, we inject non-magnetized gas with a density of $\rho_{\rm out}$, a velocity of $v_{\rm ff}(r_{\rm out})$, and a temperature of $5\times 10^{7}~{}{\rm K}$ as the fallback material.
We adopt numerical floors for density and pressure to improve numerical stability: $\rho_{\rm fl}=10^{-4}\rho_{\rm out}(r/r_{\rm PNS})^{-3}$ and $p_{\rm fl}=10^{-6}\rho_{\rm out}(\Gamma-1)(r/r_{\rm PNS})^{-4}$, respectively. In regions where the kinetic or magnetic energy is much greater than the thermal energy, small errors can lead to a negative pressure. We can avoid the problem with these floors.
4 Numerical Results
4.1 Time evolution of general structures
Figure 3: Left: radial profiles of three different pressures: the magnetic pressure $p_{\rm mag}$ (blue), the radiation pressure $p$, and the ram pressure $p_{\rm ram}$. Dashed lines indicate the magnetospheric radius $r_{\rm M}$ and the shock radius $r_{\rm sh}$. The thin line represents the radial dependence of $p$ in Equation (3). Right: various time scales against radius. We express the dynamical time scale as $t_{\rm dyn}=r/v^{(r)}$ and the cooling time scales due to the pair process as $t_{\rm pair}=e/\dot{q}_{\rm pair}$ and the URCA process as $t_{\rm URCA}=e/\dot{q}_{\rm URCA}$. The shaded region stands for a region of $r<r_{\rm M}$. Both panels are results in the fiducial model, B14Mm5R10, at $t=0.118~{}{\rm s}$.
We first present the accretion structure in our fiducial model, B14Mm5R10. Subsequently, we describe the remaining models of $r_{\rm PNS}=10~{}{\rm km}$ in comparison with the fiducial model. We show the results of $r_{\rm PNS}=20~{}{\rm km}$ in Section 4.4.
The left panel of Figure 3 presents the radial profiles of three different pressures in the fiducial model at $t=0.118~{}{\rm s}$: $p_{\rm mag}$ (blue), $p$ (gold), and $p_{\rm ram}=\rho(v^{(r)})^{2}$ (gray), where $v^{(r)}=u^{(r)}/u^{(t)}$. As mentioned in Figure 1, there are three distinct regions: the free-fall region, the post-shock region, and the PNS magnetosphere. These regions correspond to where $p_{\rm ram}$, $p$, and $p_{\rm mag}$ are the highest among the three, respectively. We define the magnetospheric radius $r_{\rm M}$ and the shock radius $r_{\rm sh}$ as the radii at which $p_{\rm mag}=p$ and $p=p_{\rm ram}$, respectively. We then obtain $r_{\rm M}=20~{}{\rm km}$ and $r_{\rm sh}=300~{}{\rm km}$ at this time (see dashed lines).
The PNS magnetic field decelerates the accreting gas, driving an accretion shock. We define the time at which the accretion shock forms as $t_{\rm shock}$. More specifically, we define $t_{\rm shock}$ as the time when a region satisfying $p>\max{(p_{\rm mag},p_{\rm ram})}$ and $p_{\rm rad}>10^{24}~{}{\rm dyn~{}cm^{-2}}$ first appears. As presented in Table 1, $t_{\rm shock}$ is almost the same for all models, indicating that $t_{\rm shock}$ is insensitive to the initial field strength of the PNS.
The thin solid line in the left panel of Figure 3 represents the radial dependence given in Equation (3). The resulting $\rho$, $p$, and $v^{(r)}$ agree with Equations (2) - (4). By comparing the thin line with the gold line, we find that the resulting pressure $p$ follows the dependence of $r^{-4}$. At $r=r_{\rm sh}$, the jump condition of $p=(6/7)p_{\rm ram}$ is satisfied. Although not shown in this panel, $\rho$ and $v^{(r)}$ are also consistent with analytical solutions (Equations 2 and 4). These analytic solutions are valid for a standing shock. Therefore, when the shock expansion speed $v_{\rm sh}=dr_{\rm sh}/dt$ is not negligible compared to $|v^{(r)}|$ at $r=r_{\rm sh}$, the resulting $\rho$, $p$, and $v^{(r)}$ deviate from the analytic solutions.
The right panel of Figure 3 shows the various time scales versus radius in the fiducial model at $t=0.118~{}{\rm s}$: the dynamical time scale $t_{\rm dyn}=r/v^{(r)}$ (blue), and cooling time scales due to the pair process $t_{\rm pair}=e/\dot{q}_{\rm pair}$ (gold) and the URCA process $t_{\rm URCA}=e/\dot{q}_{\rm URCA}$ (black). At this time, neutrino cooling is ineffective. The shaded region represents a region of $r<r_{\rm M}$. We express the time scale of the neutrino cooling as $t_{\rm cool}=\min(t_{\rm pair},t_{\rm URCA})$. For $r>r_{\rm M}$, the dynamical time scale is shorter than the cooling time scale, indicating the ineffective neutrino cooling. This panel also shows $t_{\rm pair}<t_{\rm URCA}$, which implies $\dot{q}_{\rm pair}>\dot{q}_{\rm URCA}$.
Figure 4: The time-sequenced images of $p_{\rm ram}$ (left), $p$ (middle), and $p_{\rm mag}$ (right) in the fiducial model, B14Mm5R10. Dashed lines in the left panel represent the shock formation time $t_{\rm shock}$ and the shock stalling time $t_{\rm stall}$. For $t_{\rm shock}<t<t_{\rm stall}$, the shock expands over time. However, it stalls at $t=t_{\rm stall}$ and recedes for $t>t_{\rm stall}$ due to the neutrino cooling. The white line in the right panel indicates the radius of $\beta=1$. The region on the lower and upper sides of this line correspond to $\beta<1$ and $\beta>1$, respectively.
Figure 4 displays time-sequenced images of the radial profiles of $p_{\rm ram}$ (left), $p$ (middle), and $p_{\rm mag}$ (right) in the fiducial model. For $t>t_{\rm shock}$, both $p_{\rm ram}$ and $p$ change discontinuously at $r=r_{\rm sh}$. In the early stage, $r_{\rm sh}$ expands over time. The dashed lines in the left panel indicate the times $t=t_{\rm shock}$ and $t=t_{\rm stall}$, where $t_{\rm stall}$ is the time when the expansion of $r_{\rm sh}$ stops (i.e., $v_{\rm sh}=0$). We write the ratio of $p_{\rm ram}+p$ to $p_{\rm mag}$ as $\beta$. The line in the right panel shows the radius where $\beta=1$. This radius decreases rapidly in the early time, indicating the compression of the PNS fields. The $\beta=1$ line lies in the vicinity of the PNS surface, showing the formation of the thin magnetosphere.
In the later stage ($t\geq t_{\rm stall}$), the accretion shock stalls and recedes due to the neutrino cooling. As the accreted mass accumulates above the PNS magnetosphere, the pressure just above it increases. According to the relation of $T=(3p/a_{\rm rad})^{1/4}$, $T$ at $r\sim r_{\rm M}$ also increases. Since neutrino cooling is more effective at a higher temperature, $t_{\rm cool}$ can become smaller than $t_{\rm dyn}$ at $r\sim r_{\rm M}$. The effective neutrino cooling reduces $p$ at $r\sim r_{\rm M}$, and the radiation pressure is no longer able to support gravity. For this reason, the post-shock region collapses. We will provide the dependence of the shock stalling time scale on $\dot{M}_{\rm fb}$ and $B_{\rm PNS}$ in Section 4.3.
We note that the neutrino cooling works effectively not around $r\sim r_{\rm PNS}$, but around $r\sim r_{\rm M}$. We define the pressure scale height at $r=r_{\rm M}$ as $H_{\rm M}=r_{\rm M}^{2}p/({M\rho})$. We find that $t_{\rm cool}$ is shorter than $t_{\rm dyn}$ for $r_{\rm M}\lesssim r\lesssim r_{\rm M}+H_{\rm M}$ when $t=t_{\rm stall}$. Additionally, the gas density at $r\sim r_{\rm M}$ rapidly increases for $t>t_{\rm stall}$ due to the neutrino cooling, potentially leading to the formation of a new crust that confines the PNS magnetosphere. The increase in $\rho$ for $t>t_{\rm stall}$ implies that the shock stalling time scale corresponds to the waiting time for the new crust formation. We will discuss this point in more detail in Section 4.3.
In models with a magnetized PNS and low mass-accretion rates ($B_{\rm PNS}>0$ and $\dot{M}_{\rm fb}=10^{-(5-4)}~{}{\rm M}_{\odot}~{}{\rm s}^{-1}$), the results are very similar to those of the fiducial model. However, in models with high mass-accretion rates ($\dot{M}_{\rm fb}=10^{-(3-2)}~{}{\rm M}_{\odot}~{}{\rm s}^{-1}$) or in non-magnetized PNS models ($B_{\rm PNS}=0$), the results differ from those of the fiducial model.
In models with high mass accretion rates, neutrino cooling is effective even at $t\sim t_{\rm shock}$. A higher $\dot{M}_{\rm fb}$ leads to larger $\rho$ and $T$ in the post-shock region, resulting in a greater $\dot{q}$. As a result, $t_{\rm cool}$ can be shorter than $t_{\rm dyn}$. We also find $t_{\rm URCA}\sim t_{\rm pair}$ in these models, which implies that both the URCA and pair processes are equally important. Such effective neutrino cooling leads to a reduction in the effective adiabatic index of the gas. Therefore, the resulting $\rho$, $p$, and $v^{(r)}$ are slightly steeper than $\propto r^{-3}$, $\propto r^{-4}$, and $\propto r$, respectively.
In the non-magnetized cases, the post-shock region is formed in $r_{\rm PNS}<r<r_{\rm sh}$, while the free-fall region is located in $r>r_{\rm sh}$. The cooling time scale is shorter than the dynamical time scale for $r_{\rm PNS}\lesssim r\lesssim r_{\rm PNS}+H$ at $t=t_{\rm stall}$, which is consistent with the assumption in Chevalier (1989) (see Section 2).
We had to stop our calculation around $t\approx 1.2t_{\rm stall}$ because the pressure scale height $H_{\rm M}$ or $H$ became smaller than $\Delta r$ due to a significant increase in $\rho$. This increase in $\rho$ makes it impossible to resolve the stratified atmosphere.
4.2 Magnetospheric and shock radii
Figure 5: Top and middle: plots for the magnetospheric (top) and shock radii (middle) as a function of the fallback accretion rate. All plots are the results when $t=t_{\rm stall}$. The solid lines indicate the semi-analytical solutions calculated from Equations (16), (18), and (19). The shaded regions in the top and middle panels denote where the deviation from the semi-analytic solution for $B_{\rm PNS}=10^{15}~{}{\rm G}$ is within 20% and 50%, respectively. We also plot the analytical solution of $r_{\rm sh}$ for $B_{\rm PNS}=0$ (Equation 8). Bottom: shock stalling time scale ($t^{\prime}_{\rm stall}=t_{\rm stall}-t_{\rm shock}$) against the fallback accretion rate. Solid lines represent the fitting results. All panels are the results of $r_{\rm PNS}=10~{}{\rm km}$.
We show the dependence of $r_{\rm M}$ and $r_{\rm sh}$ at $t=t_{\rm stall}$ on $\dot{M}_{\rm fb}$ and $B_{\rm PNS}$. Additionally, based on semi-analytical models, we demonstrate the following two points: (1) the magnetospheric radius is determined by the pressure balance between $p$ and $p_{\rm mag}$; (2) the shock radius is obtained from the energy balance at $r=r_{\rm M}$, which describes that the released gravitational energy is lost through the neutrino cooling.
The top panel of Figure 5 shows the parameter dependence of $r_{\rm M}$ at $t=t_{\rm stall}$. As expected from the pressure balance, the panel demonstrates that $r_{\rm M}$ is a decreasing function of $\dot{M}_{\rm fb}$ but an increasing function of $B_{\rm PNS}$.
The middle panel of Figure 5 presents the dependence of $r_{\rm sh}$ at $t=t_{\rm stall}$ on $\dot{M}_{\rm fb}$ and $B_{\rm PNS}$. We also plot $r_{\rm sh}$ in models of $B_{\rm PNS}=0$ for comparison. The plots show that $r_{\rm sh}$ also decreases with $\dot{M}_{\rm fb}$ and increases with $B_{\rm PNS}$, although the dependence on $B_{\rm PNS}$ is weaker than that of $r_{\rm M}$. When $r_{\rm M}$ is smaller, $\rho$ and $T$ just above it become higher due to stronger gravitational acceleration. Given the parameter dependence of the neutrino cooling (see Equations 6 and 7), the cooling becomes more effective, resulting in a smaller $r_{\rm sh}$. A comparison with the non-magnetized case ($B_{\rm PNS}=0$) indicates that the PNS magnetic field significantly affects $r_{\rm sh}$ only when $B_{\rm PNS}\gtrsim 10^{14}~{}{\rm G}$, within the range of $\dot{M}_{\rm fb}$ investigated here.
[…]
$p_{\rm mag}=\frac{\bar{B}^{2}}{8\pi}=\frac{\mu^{2}}{2\pi}\frac{1}{\left(r_{\rm M}^{2}-r_{\rm PNS}^{2}\right)^{2}}\left(\frac{1}{r_{\rm PNS}}\right)^{2},$ (15)
where $\mu=B_{\rm PNS}r_{\rm PNS}^{3}$ is the PNS magnetic moment. Substituting Equations (3) and (15) into the pressure balance equation $p=p_{\rm mag}$, we get the following equation that relates $r_{\rm M}$ and $r_{\rm sh}$:
$\left[\left(\frac{\mu^{2}}{r_{\rm PNS}}\right)^{2}-\frac{3}{7}\left(\frac{2r_{\rm g}}{r_{\rm sh}}\right)^{1/2}\dot{M}_{\rm fb}r_{\rm sh}^{2}\right]r_{\rm M}^{4}+\frac{6}{7}\left(\frac{2r_{\rm g}}{r_{\rm sh}}\right)^{1/2}\dot{M}_{\rm fb}r_{\rm sh}^{2}r_{\rm PNS}^{2}r_{\rm M}^{2}-\frac{3}{7}\left(\frac{2r_{\rm g}}{r_{\rm sh}}\right)^{1/2}\dot{M}_{\rm fb}r_{\rm sh}^{2}r_{\rm PNS}^{4}=0.$ (16)
Since the neutrino cooling is effective mainly at $r\sim r_{\rm M}$, we adopt the following energy equation instead of Equation (5):
$\frac{M_{\rm PNS}\dot{M}_{\rm fb}}{r_{\rm M}}=4\pi r_{\rm M}^{2}H_{\rm M}\dot{q}.$ (17)
Substituting $\dot{q}_{\rm pair}$ and $\dot{q}_{\rm URCA}$ into Equation (17), we obtain the following shock radii, respectively (c.f., Equation 8 and 9):
$r_{\rm sh,pair}$ $=$ $9.5\times 10^{2}~{}[{\rm km}]\left(\frac{r_{\rm M}}{20~{}{\rm km}}\right)^{40/27}\left(\frac{M_{\rm PNS}}{1.4~{}{\rm M_{\odot}}}\right)^{-1/27}\left(\frac{\dot{M}_{\rm fb}}{10^{-5}~{}{\rm M_{\odot}~{}s^{-1}}}\right)^{-10/27},$ (18)
$r_{\rm sh,URCA}$ $=$ $1.1\times 10^{3}~{}[{\rm km}]\left(\frac{r_{\rm M}}{20~{}{\rm km}}\right)^{4/3}\left(\frac{M_{\rm PNS}}{1.4~{}{\rm M_{\odot}}}\right)^{1/5}\left(\frac{\dot{M}_{\rm fb}}{10^{-5}~{}{\rm M_{\odot}~{}s^{-1}}}\right)^{-2/5}.$ (19)
For given parameters of $\dot{M}_{\rm fb}$ and $B_{\rm PNS}$, we numerically solve Equations (16), (18), and (19) for $r_{\rm M}$ and $r_{\rm sh}$. We solve these equations with a double iteration scheme. We adopt $r_{\rm M,init}=10^{3}~{}{\rm km}$ as the initial guess of $r_{\rm M}$. The steps of the iterative scheme are as follows: (i) Using $r_{\rm M,init}$, we compute the shock radius from $r_{\rm sh}=\min(r_{\rm sh,pair},r_{\rm sh,URCA})$. (ii) Based on $r_{\rm sh}$, we solve Equation (16) for $r_{\rm M}$ using the Newton-Raphson method. We check the convergence of $\log(r_{\rm M})$. If a relative error of $\log(r_{\rm M})$ falls below $10^{-5}$, then we proceed to the next step. (iii) If the obtained $r_{\rm M}$ does not coincide with the given $r_{\rm M,init}$, then we set the next guess of $r_{\rm M,init}$ and return to step (i). The next guess is estimated from $\log(r_{\rm M,init})-10^{-6}\log(r_{\rm M,init})$. We continue the iteration procedure until the relative accuracy $|\log(r_{\rm M,init})-\log(r_{\rm M})|/|\log(r_{\rm M})|$ reaches $0.1\%$.
The solid lines in the top and middle panels of Figure 5 stand for the semi-analytical solutions. We also plot the solution of $r_{\rm sh}$ for $B_{\rm PNS}=0$ (Equation 8) with a black line. The shaded regions in the top and middle panels denote where the deviations from the semi-analytic solutions for $B_{\rm PNS}=10^{15}~{}{\rm G}$ remain within 20 % and 50 %, respectively. These two panels show that our semi-analytical solutions are consistent with the numerical results of $r_{\rm M}$ and $r_{\rm sh}$ within the deviations of $\approx$ 20% and $\approx$ 50%, respectively. We get the same results for $B_{\rm PNS}=10^{13}~{}{\rm G}$ and $10^{14}~{}{\rm G}$.
4.3 Shock stalling time scale, $t^{\prime}_{\rm stall}$
As described in Section 4.1, the density just above the magnetosphere begins to increase rapidly at $t=t_{\rm stall}$ due to neutrino cooling. This abrupt increase can potentially lead to the formation of the new crust that confines the PNS magnetosphere. As a measure of the waiting time for the new crust formation, we define the shock stalling time scale, $t^{\prime}_{\rm stall}=t_{\rm stall}-t_{\rm shock}$, and investigate its parameter dependence. Additionally, we aim to obtain a fitting formula of $t^{\prime}_{\rm stall}$ to discuss a necessary condition for the submergence of the PNS field.
We first focus on the non-magnetized case as a reference (see black circles in the bottom panel of Figure 5). We analyze the results to find the following: when the accretion rate is high ($\dot{M}_{\rm fb}=10^{-(3-2)}~{}{\rm M_{\odot}~{}s^{-1}}$), the shock stalling time scale is characterized by the cooling time scale, namely $t^{\prime}_{\rm stall}\approx t_{\rm cool}$. However, when the accretion rate is low ($\dot{M}_{\rm fb}=10^{-(5-4)}~{}{\rm M_{\odot}~{}s^{-1}}$), $t^{\prime}_{\rm stall}$ is significantly longer than $t_{\rm cool}$ due to the time delay in the onset of effective neutrino cooling. We have confirmed that, in all models, the sound-crossing timescale across the region between $r=r_{\rm M}$ and $r=r_{\rm sh}$, $t_{\rm sound}$, is shorter than $t_{\rm cool}$ for $t\leq t_{\rm stall}$. The relation $t_{\rm sound}<t_{\rm cool}$ indicates that hydrostatic equilibrium is almost achieved in the post-shock region.
The black solid line in the bottom panel of Figure 5 denotes the fitting function for the results of the non-magnetized models. The functional form is
$t^{\prime}_{\rm stall,non}(\dot{M}_{\rm fb})=9.8\times 10^{-2}~{}[{\rm s}]~{}\left(\frac{\dot{M}_{\rm fb}}{{10^{-5}~{}\rm M_{\odot}~{}s^{-1}}}\right)^{-0.69}.$ (20)
Since $\dot{q}$ is large for a high $\dot{M}_{\rm fb}$, $t^{\prime}_{\rm stall}$ decreases with $\dot{M}_{\rm fb}$.
The bottom panel of Figure 5 indicates that $t^{\prime}_{\rm stall}$ gets longer for a stronger $B_{\rm PNS}$, as the neutrino cooling becomes more ineffective (see Section 4.2). To derive a fitting function for the magnetized models, we hypothesize that $t^{\prime}_{\rm stall}$ for the magnetized cases is the sum of $t^{\prime}_{\rm stall,non}$ and an additional term that depends on $\dot{M}_{\rm fb}$ and $B_{\rm PNS}$, that is,
$t^{\prime}_{\rm stall}(\dot{M}_{\rm fb},B_{\rm PNS})=t^{\prime}_{\rm stall,non}(\dot{M}_{\rm fb})+f(\dot{M}_{\rm fb},B_{\rm PNS}),$ (21)
where $f(\dot{M}_{\rm fb},B_{\rm PNS})$ denotes the additional term. We then derive the fitting function for $f(\dot{M}_{\rm fb},B_{\rm PNS})$ based on the magnetized models. In models B13Mm2R10 and B14Mm2R10, $t^{\prime}_{\rm stall}$ is slightly shorter than $t^{\prime}_{\rm stall,non}$ because $r_{\rm M}\approx r_{\rm PNS}$. Therefore, we exclude $t^{\prime}_{\rm stall}$ of these models from this fitting.
The result is as follows:
$f(\dot{M}_{\rm fb},B_{\rm PNS})=2.0\times 10^{-1}~{}[{\rm s}]~{}\left(\frac{\dot{M}_{\rm fb}}{{10^{-5}~{}\rm M_{\odot}~{}s^{-1}}}\right)^{-1.2}\left(\frac{B_{\rm PNS}}{{10^{14}~{}\rm G}}\right)^{1.3}.$ (22)
The resulting fitting functions are plotted as solid lines in the bottom panel of Figure 5. It turns out that the accuracy of the fitting formula is better than $\approx 96\%$.
4.4 The dependence on the PNS radius
The results described above are for $r_{\rm PNS}=10~{}{\rm km}$. To clarify how the dynamics of the supernova fallback depend on $r_{\rm PNS}$, we study the case for $r_{\rm PNS}=20~{}{\rm km}$ (see Table 1).
The dependence of $r_{\rm M}$ and $r_{\rm sh}$ is as follows. As the gravitational acceleration at the PNS surface becomes smaller with increasing $r_{\rm PNS}$, both the radiation pressure and the cooling rate decrease. Consequently, both $r_{\rm M}$ and $r_{\rm sh}$ for $r_{\rm PNS}=20~{}{\rm km}$ are larger than those for $r_{\rm PNS}=10~{}{\rm km}$. We confirm that our semi-analytic models are also consistent with these radii for $r_{\rm PNS}=20~{}{\rm km}$ (see Section 4.2).
A larger $r_{\rm PNS}$ leads to a longer $t^{\prime}_{\rm stall}$. We derive the following functional form in the same manner as in Section 4.3:
$t^{\prime}_{\rm stall}(\dot{M}_{\rm fb},B_{\rm PNS})=1.5~{}[{\rm s}]~{}\left(\frac{\dot{M}_{\rm fb}}{{10^{-5}~{}\rm M_{\odot}~{}s^{-1}}}\right)^{-0.83}+1.2\times 10^{1}~{}[{\rm s}]~{}\left(\frac{\dot{M}_{\rm fb}}{{10^{-5}~{}\rm M_{\odot}~{}s^{-1}}}\right)^{-1.3}\left(\frac{B_{\rm PNS}}{{10^{14}~{}\rm G}}\right)^{1.2},$ (23)
The accuracy of this fitting formula is better than $\approx 94\%$.
4.5 Necessary condition for the submergence of the PNS magnetic field
We consider that the submergence of the PNS fields requires the abrupt increase in the gas density in the post-shock region. The abrupt increase occurs when a sufficient amount of accreted matter accumulates at $r\sim r_{\rm M}$ and the gas density and temperature become high enough to induce rapid cooling (that is why the shock starts to stall at that time; see Section 4.1). This may not be true when the supernova fallback ceases before the cooling becomes effective. As we have ignored the duration of the supernova fallback in the above analysis, we investigate this effect.
To investigate the impact of the finite duration of the fallback, we perform a test simulation based on our fiducial model, B14Mm5R10. The setup of the test simulation is as follows. As an initial condition, we take the physical quantities from the fiducial model at $t\approx 0.07~{}{\rm s}$. We note $t_{\rm stall}\approx 0.25~{}{\rm s}$ in the fiducial model (see the dashed line in Figure 4). By setting $\rho=\rho_{\rm fl}$ and $p=p_{\rm fl}$ at $r=r_{\rm out}$, we effectively stop the fallback accretion at $t\sim 0.1~{}{\rm s}$. In ghost cells outside $r=r_{\rm out}$, we impose $u^{r}=\max(u^{r},0)$ and adopt the zero-gradient boundary condition for $\rho$ and $p$.
Figure 6: Time-sequenced images illustrating the radial profiles of the three different pressures in a test simulation. In this test simulation, the initial conditions are extracted from the physical quantities at $t\approx 0.07~{}{\rm s}$ in model B14Mm5R10. By adopting $\rho=\rho_{\rm fl}$ and $p=p_{\rm fl}$ at $r=r_{\rm out}$, we effectively stop the fallback accretion at $t\sim 0.1~{}{\rm s}$ (see Section 4.5 for details). All labels are the same as in Figure 4.
Figure 6 displays time-sequenced images of radial profiles of the three different pressures. Due to the absence of the free-fall region, the post-shock gas expands to reduce its density and temperature (see the middle panel). The simulation does not show the abrupt increase in the density, suggesting that the PNS magnetosphere will not be submerged for $t^{\prime}_{\rm stall}>t_{\rm fb}$. Therefore, the condition $t^{\prime}_{\rm stall}<t_{\rm fb}$ needs to be satisfied for the submergence.
5 Discussion
5.1 Implications for the Diversity in Young isolated NS
We propose possible solutions for the supernova fallback based on the following two conditions and discuss the diversity of young isolated NSs. The first condition is a necessary condition for the submergence of the PNS field (Section 4.5). The second condition is related to an effect of the PNS rotation, which is ignored in our numerical models. The rotating PNS can drive relativistic outflows and repel the fallback matter (e.g., Zhong et al., 2021). We take this effect into account based on the theoretical study of Zhong et al. (2021).
We have suggested that the inequality $t^{\prime}_{\rm stall}<t_{\rm fb}$ needs to be satisfied for the submergence of the PNS field (see Section 4.5). Using Equations (21) and (23), we can rewrite the relation $t^{\prime}_{\rm stall}<t_{\rm fb}$ as
$9.8\times 10^{-2}~{}[{\rm s}]~{}\left(\frac{\dot{M}_{\rm fb}}{{10^{-5}~{}\rm M_{\odot}~{}s^{-1}}}\right)^{-0.69}+2.0\times 10^{-1}~{}[{\rm s}]~{}\left(\frac{\dot{M}_{\rm fb}}{{10^{-5}~{}\rm M_{\odot}~{}s^{-1}}}\right)^{-1.2}\left(\frac{B_{\rm PNS}}{{10^{14}~{}\rm G}}\right)^{1.3}$ $<$ $t_{\rm fb},$ (24)
$1.5~{}[{\rm s}]~{}\left(\frac{\dot{M}_{\rm fb}}{{10^{-5}~{}\rm M_{\odot}~{}s^{-1}}}\right)^{-0.83}+1.2\times 10^{1}~{}[{\rm s}]~{}\left(\frac{\dot{M}_{\rm fb}}{{10^{-5}~{}\rm M_{\odot}~{}s^{-1}}}\right)^{-1.3}\left(\frac{B_{\rm PNS}}{{10^{14}~{}\rm G}}\right)^{1.2}$ $<$ $t_{\rm fb},$ (25)
for $r_{\rm PNS}=10~{}{\rm km}$ and $r_{\rm PNS}=20~{}{\rm km}$, respectively.
The condition related to the PNS rotation is as follows (Zhong et al., 2021). The spin-down luminosity is expressed as (Spitkovsky, 2006)
$L_{\rm spin}=\frac{B_{\rm PNS}^{2}r_{\rm PNS}^{6}\Omega_{\rm PNS}^{4}}{4}(1+\sin^{2}\chi),$ (26)
where $\Omega_{\rm PNS}=2\pi/P$ is the angular velocity of the PNS, $P$ is the spin period, and $\chi$ is the offset angle between the rotation and dipole axes. We take $\chi=0$ for simplicity. Zhong et al. (2021) assumed that the relativistic outflows and fallback matter typically encounters at the fallback radius, $r_{\rm fb}\sim(M_{\rm PNS}t_{\rm fb}^{2})^{1/3}$. From the condition of $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}(r_{\rm fb})$, we obtain (see also Shigeyama & Kashiyama, 2018, for more generalized case)
$\dot{M}_{\rm fb}>\frac{1}{2}L_{\rm spin}M_{\rm PNS}^{-2/3}t_{\rm fb}^{2/3}.$ (27)
Since the spin-down time scale, $\sim 10~{}{\rm yr}$, is much longer than $t_{\rm fb}$, the spin-down luminosity $L_{\rm spin}$ is regarded as a constant with time. We employ $P=0.02~{}{\rm s}$ as a fiducial value in Equation (27), which is shorter than the pulse periods observed in CCOs ($\sim 0.1~{}{\rm s}$, Enoto et al., 2019; Borghese & Esposito, 2023).
Figure 7: Possible solutions of the supernova fallback illustrated as a function of fallback mass and time. We employ $B_{\rm PNS}=10^{13}~{}{\rm G}$ and $P=0.02~{}{\rm s}$. We show two cases of $r_{\rm PNS}=10~{}{\rm km}$ (left) and $20~{}{\rm km}$ (right). The thick and thin lines stand for $t^{\prime}_{\rm stall}=t_{\rm fb}$ (Equation 24) and $L_{\rm spin}=\dot{M}_{\rm fb}v^{2}_{\rm ff}$ (Equation 27), respectively. The fallback material submerges the PNS magnetosphere for $t^{\prime}_{\rm stall}<t_{\rm fb}$ and $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (red-shaded region). The fallback material is repelled by the spin-down power for $L_{\rm spin}>\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (blue-shaded region). The fallback material disturbs the PNS magnetosphere for $t^{\prime}_{\rm stall}>t_{\rm fb}$ and $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (orange-shaded region). We represent the critical point branching into the three different solutions as a black point. We consider that CCOs form in the red-shaded region, whereas the other types of NSs are born in the orange- and/or blue-shaded regions.
Figure 7 displays the possible solutions for $B_{\rm PNS}=10^{13}~{}{\rm G}$. Here, we have used the initial value of the accretion rate for the fallback matter, $\dot{M}_{\rm fb}=(2/3)(M_{\rm fb}/t_{\rm fb})$ (see Equation 1). The thick and thin lines represent $t^{\prime}_{\rm stall}=t_{\rm fb}$ and $L_{\rm spin}=\dot{M}_{\rm fb}v^{2}_{\rm ff}$, respectively. In the region where $t^{\prime}_{\rm stall}<t_{\rm fb}$ and $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (red-shaded region), the supernova fallback can form a new crust and submerge the PNS magnetosphere. In the region where $L_{\rm spin}>\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (blue-shaded region), the fallback material is repelled by the dipole spin-down power and therefore does not interact with the PNS magnetosphere. In this case, the NS magnetic field is determined solely by internal amplifications during the core-collapse and PNS phases. In the region where $t^{\prime}_{\rm stall}>t_{\rm fb}$ and $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (orange-shaded region), the fallback accretion is halted before the PNS magnetosphere is submerged, allowing the magnetosphere to survive. During the fallback phase, the accreting material interacts with the PNS and disturbs its magnetosphere. We speculate that the disturbance leads to an external amplification of the magnetic field. We will discuss the amplification mechanisms induced by the supernova fallback later.
Figure 7 shows that the critical points, which branch into three different solutions, are located around $(t_{\rm fb},~{}M_{\rm fb})\approx(2\times 10^{3}~{}{\rm s},~{}2\times 10^{-6}~{}{\rm M}_{\odot})$ for $r_{\rm PNS}=10~{}{\rm km}$ and $(t_{\rm fb},~{}M_{\rm fb})\approx(2\times 10^{3}~{}{\rm s},~{}10^{-4}~{}{\rm M}_{\odot})$ for $r_{\rm PNS}=20~{}{\rm km}$ (see black points). As $r_{\rm PNS}$ becomes large, a higher $M_{\rm fb}$ is required for the submergence. Although not shown in this figure, the red-shaded region becomes narrower for a stronger $B_{\rm PNS}$ within the plotted parameter range.
The solutions in Figure 7 will provide insight into the diversity of the young isolated NSs. The CCOs can form in the red-shaded region, while the other types of the NSs are born in the orange- and/or blue-shaded regions. The solutions mean that when $B_{\rm PNS}=10^{13}~{}{\rm G}$ and $P=0.02~{}{\rm s}$, the NS population can be dominated by the CCOs within the typical ranges of $M_{\rm fb}=10^{-(4-1)}~{}{\rm M}_{\odot}$ and $t_{\rm fb}=10^{0-3}~{}{\rm s}$. This may suggest that the CCOs are common in the young isolated NSs. In fact, despite the small number of the known CCOs, observational studies implied that the birth rate of the CCOs is as large as the other types of the NSs (Gaensler et al., 2000).
We note that multidimensional effects, many of which are ignored in our simulations, will modify the solutions in Figure 7. Recent multidimensional simulations of the supernova explosion suggested that the fallback gas may have a large angular momentum (see Janka et al., 2022, and references therein). When such material accretes onto the PNS, an accretion disk will form around the PNS magnetosphere (Takahashi & Ohsuga, 2017; Inoue et al., 2023, 2024). The differential rotation of the disk opens up the PNS dipole field, which enhances $L_{\rm spin}$ (Parfrey & Tchekhovskoy, 2017). Furthermore, MHD instabilities are also important. The surface at $r_{\rm M}$, where the magnetic pressure balances the radiation pressure, can be unstable to the magnetic Rayleigh-Taylor instability (see e.g., Stone & Gardiner, 2007). This instability leads to dynamic mixing around $r\sim r_{\rm M}$, suppressing the submergence of the PNS magnetosphere (Bernal et al., 2013). Therefore, the solutions require further updates. However, we consider that our results will offer valuable guidance for investigating the diversity of young isolated NSs through multidimensional simulations.
The supernova fallback may lead to an external amplification of the PNS magnetic field. When the fallback matter reaches the PNS without submerging the surface magnetic fields (orange-shaded region in Figure 7), it will chaotically disturb the magnetosphere. This chaotic process can lead to stronger and compressed multipolar magnetic fields (see discussion in Zhong et al., 2021). Furthermore, Barrère et al. (2022) proposed a magnetar formation scenario through the supernova fallback, in which the PNS field is amplified due to the Tayler-Spruit dynamo triggered by the PNS spin-up from the supernova fallback (see also Barrère et al., 2025). Although a detailed investigation of magnetic field evolution is beyond the scope of this study, we speculate that supernova fallback serves as an additional factor contributing to magnetic field amplification.
We discuss how the crystallization of the fallback material affects Figure 7. We have derived Equations (24) and (25) by comparing $t_{\rm stall}^{\prime}$ with $t_{\rm fb}$. However, the submergence of the PNS field requires the new crust formation (see Figure 1), which means that we need to take into account the time interval from when the shock stalls to when the gas density and temperature satisfy the crystallizing condition. To estimate such a time interval, we have performed an additional simulation based on model BnMm2R10 (see Appendix A for details). From this simulation, we find that the time interval is of the same order as the shock stalling time scale. Therefore, we conclude that the crystallization of the fallback material does not significantly affect Figure 7. The crystallization time scale may become longer if we work with a more complete and realistic equation of state, as the degenerate pressure suppresses the increase in the gas density. Such simulations are left for important future work (Bernal et al., 2010, 2013).
5.2 Estimation of the effects of neutrino scattering and absorption
Our numerical models assume that neutrino cooling operates in an optically thin regime (see Equation 14). To evaluate the validity of this approximation, we estimate the radii of the neutrinospheres for $\nu_{\rm e}$ and $\bar{\nu}_{\rm e}$, at which the neutrino depths $\tau_{\nu_{\rm e}}$ and $\tau_{\bar{\nu}_{\rm e}}$ become unity, respectively.
We consider neutrino scattering and absorption by free protons and neutrons. We ignore the effect of neutrino-electron scattering because its cross section is much smaller than those of neutrino-proton and neutrino-neutron scatterings (Shapiro & Teukolsky, 1983). We write the electron, proton, and neutron masses as $m_{\rm e}$, $m_{\rm p}$, and $m_{\rm n}$, respectively ($m_{\rm p}\approx m_{\rm n}\gg m_{\rm e}$). We define the cross sections for the absorption processes of $\bar{\nu}_{e}+p\rightarrow e^{+}+n$ and $\nu_{e}+n\rightarrow e^{-}+p$ as
$\sigma_{\rm abs}(\bar{\nu}_{\rm e}p)$ $=$ $\sigma_{0}\left(\frac{1+3g_{\rm A}^{2}}{4}\right)\left(\frac{\varepsilon_{\bar{\nu}_{\rm e}}-\Delta_{\rm np}}{m_{\rm e}}\right)^{2}\left[1-\left(\frac{m_{\rm e}}{\varepsilon_{\bar{\nu}_{\rm e}}-\Delta_{\rm np}}\right)^{2}\right]^{1/2},$ (28)
$\sigma_{\rm abs}(\nu_{\rm e}n)$ $=$ $\sigma_{0}\left(\frac{1+3g_{\rm A}^{2}}{4}\right)\left(\frac{\varepsilon_{\nu_{\rm e}}+\Delta_{\rm np}}{m_{\rm e}}\right)^{2}\left[1-\left(\frac{m_{\rm e}}{\varepsilon_{\nu_{\rm e}}+\Delta_{\rm np}}\right)^{2}\right]^{1/2},$ (29)
respectively (Burrows et al., 2006). Here, $\sigma_{0}=1.705\times 10^{-44}~{}{\rm cm^{2}}$, $g_{\rm A}$ is the axial–vector coupling constant ($g_{\rm A}\approx-1.23$), $\varepsilon_{\nu_{\rm e}}$ is the energy of $\nu_{\rm e}$, $\varepsilon_{\bar{\nu}_{\rm e}}$ is the energy of $\bar{\nu}_{\rm e}$, and $\Delta_{\rm np}=m_{\rm n}-m_{\rm p}$. We assume that the correction for weak magnetism and recoil is negligible for simplicity. We describe the cross sections in the elastic limit of neutrino-proton and neutrino-neutron scatterings as follows (see e.g., Shapiro & Teukolsky, 1983; Suwa et al., 2019):
$\sigma_{\rm sca,p}\sim\sigma_{\rm sca,n}=\frac{\sigma_{0}}{4}\left(\frac{\varepsilon_{\nu}}{m_{\rm e}}\right)^{2}.$ (30)
The method for estimating the radii of the neutrinospheres is as follows. We denote the number densities of electrons, protons, and neutrons as $n_{\rm e}$, $n_{\rm p}$, and $n_{\rm n}$, respectively. We assume the charge neutrality condition $n_{\rm e}=n_{\rm p}$. We can then write the mass density as $\rho=m_{\rm p}(n_{\rm p}+n_{\rm n})$. We represent the electron fraction as $Y_{\rm e}=n_{\rm p}/(n_{\rm p}+n_{\rm n})$. From these definitions, we get the following relations: $n_{\rm p}=\rho Y_{\rm e}/{m_{\rm p}}$ and $n_{\rm n}={(1-Y_{\rm e})}\rho/m_{\rm p}$. The detailed estimation of the value of $Y_{e}$ is beyond the scope of this study, and we adopt $Y_{\rm e}=0.5$ as a fiducial value for simplicity (Fryer et al., 2006). From the definitions explained above, we calculate the radii of neutrinospheres using the following equations:
[…]
Since the typical gas temperature at $r\sim r_{\rm M}$ is $T\sim 10^{11}~{}{\rm K}$, we calculate $\tau_{\nu_{\rm e}}$ and $\tau_{\bar{\nu}_{\rm e}}$ for $\varepsilon_{\nu_{\rm e}}=\varepsilon_{\bar{\nu}_{\rm e}}=10~{}{\rm MeV}$ (see also Akaho et al., 2024).
In all models, when $t=t_{\rm stall}$, both $\tau_{\nu_{\rm e}}$ and $\tau_{\bar{\nu}_{\rm e}}$ are much smaller than unity even at $r=r_{\rm M}$, suggesting that neutrino scattering and absorption by free protons and neutrons can be neglected (see Sumiyoshi & Yamada, 2012, for more detailed estimation).
6 Summary
To understand the observational diversity of the young isolated NSs, we performed one-dimensional GRMHD simulations of supernova fallback onto a magnetized PNS, while taking neutrino cooling into account. We demonstrated that the PNS magnetosphere, the post-shock region, and free-fall region appear (Figure 3). The accretion shock expands in an early stage, whereas it stalls and recedes in a later stage due to the neutrino cooling (Figure 4). Based on our numerical models, we derived a necessary condition for the submergence of the PNS magnetic field. We summarize our findings in the following.
The one-dimensional simulations enabled us to study the time evolution of the fallback material over wide parameter ranges including the fallback mass accretion rate ($\dot{M}_{\rm fb}=10^{-(5-2)}~{}{\rm M}_{\odot}~{}{\rm s}^{-1}$), the magnetic field strength at the PNS surface ($B_{\rm PNS}=10^{13-15}~{}{\rm G}$, and $B_{\rm PNS}=0$ for comparison), and the PNS radius ($r_{\rm PNS}=10$ and $20~{}{\rm km}$). From these simulations, we revealed that an abrupt increase in the gas density occurs after the shock stalls because of the neutrino cooling. The abrupt increase potentially leads to the formation of a new crust that confines the PNS magnetosphere.
We developed semi-analytic models that are consistent with the magnetospheric and shock radii ($r_{\rm M}$ and $r_{\rm sh}$, respectively) when the shock stalls (Section 4.2). The semi-analytic models are based on the following two relations: (1) the pressure balance between the magnetic pressure of the PNS field and the radiation pressure in the post-shock region (Equation 16), and (2) the balance between the release rate of the gravitational energy and energy loss rate due to neutrino cooling at $r=r_{\rm M}$ (Equation 17). Therefore, both $r_{\rm M}$ and $r_{\rm sh}$ decrease with $\dot{M}_{\rm fb}$ and increase with $B_{\rm PNS}$ (see top and middle panels of Figure 5). Our semi-analytic model reduces to the analytical solution in Chevalier (1989) when $B_{\rm PNS}=0$.
We found that the shock stalling time scale can be regarded as the waiting time for the new crust formation (Section 4.3). When the accretion rate is high ($\dot{M}_{\rm fb}=10^{-(3-2)}~{}{\rm M_{\odot}~{}s^{-1}}$), the shock stalling time scale is nearly equal to the time scale of the neutrino cooling. When the accretion rate is low ($\dot{M}_{\rm fb}=10^{-(5-4)}~{}{\rm M_{\odot}~{}s^{-1}}$), the shock stalling time scale consists of the cooling time scale and the time delay for neutrino cooling to become effective. Thus, the shock stalling time scale is long for a low $\dot{M}_{\rm fb}$ or a strong $B_{\rm PNS}$ (see the bottom panel of Figure 5).
The dynamics of the supernova fallback depends on $r_{\rm PNS}$ (Section 4.4). Since the gravitational acceleration at the PNS surface is smaller for larger $r_{\rm PNS}$, the radiation pressure and the cooling rate are reduced. Therefore, both $r_{\rm M}$ and $r_{\rm sh}$ are larger for larger $r_{\rm PNS}$. Furthermore, due to the lower cooling rate, the shock stalling time scale becomes longer for larger $r_{\rm PNS}$.
We proposed a necessary condition for the submergence of the PNS magnetosphere (Section 4.5). The submergence requires the abrupt increase in the gas density in the post-shock region during the fallback phase. Therefore, the shock stalling time scale needs to be shorter than the fallback time scale for the submergence.
Based on the necessary condition, we updated possible solutions of the supernova fallback to better understand the origin of the diversity of NS magnetic fields (Figure 7 in Section 5.1). The solutions require further updates because they ignore some multidimensional effects. However, we believe that the presented scheme will be a new benchmark for the scenarios of the NS diversity.
We thank Ryuichiro Akaho, Yudai Suwa, and Kenta Dohi for fruitful discussions. This work was supported by JSPS KAKENHI grant Nos. JP24KJ0143, JP25K17439 (A.I.), JP22K14074, JP22KK0043, JP21H04487 (S.T.), JP24K00668, JP23H04899, JP22H00130 (K.K.), and JP24K00672, JP21H04488, JP24K00678 (H.R.T). A part of this research has been funded by the MEXT as ”Program for Promoting Researches on the Supercomputer Fugaku” (Toward a unified view of the universe: from large-scale structures to planets, JPMXP1020200109; A.I. and H.R.T.). Numerical computations were performed with computational resources provided by Cray XC 50 at the Center for Computational Astrophysics (CfCA) of the National Astronomical Observatory of Japan (NAOJ), the FUJITSU Supercomputer PRIMEHPC FX1000 and FUJITSU Server PRIMERGY GX2570 (Wisteria/BDEC-01) at the Information Technology Center, The University of Tokyo.
Appendix A Time scale for the crust formation
Our simulations demonstrated that neutrino cooling effectively works on a time scale of $t^{\prime}_{\rm stall}$, leading to an abrupt increase in the gas density at $r\approx r_{\rm M}$. As a result, a new crust will form just above the PNS magnetosphere. We aim to estimate the time scale for the crust formation. We define the time scale as the period between $t_{\rm stall}$ and the time when the crystallization begins.
When Coulomb energy dominates the thermal energy in the accreted matter, ions will crystallize into a periodic lattice structure to form the NS crust. The crystallization of ions occurs if the gas temperature cools down to a melting temperature $T_{\rm m}=(Z^{2}e_{\rm c}^{2})(\Gamma_{\rm m}k)^{-1}(4\pi n_{\rm ion}/3)^{1/3}$ (see e.g., Shapiro & Teukolsky, 1983; Chamel & Haensel, 2008), where $Z$ is the typical proton number of the lattice component, $e_{\rm c}$ is the elementary charge, $\Gamma_{\rm m}$ is the ratio of Coulomb to thermal energies of the lattice at the melting point, $k$ is the Boltzmann constant, and $n_{\rm ion}$ is the number density of ion. For simplicity, we assume that the crust is formed of a perfect crystal with a single nuclear species at lattice sites (Chamel & Haensel, 2008). Its mass number and mass fraction are denoted by $A=Z/Y_{\rm e}$ and $x_{\rm a}$, respectively. We then write the number density of ions as $n_{\rm ion}=(\rho x_{\rm a})/(Am_{\rm p})$. From these definitions, we can rewrite the melting temperature as (Suwa, 2014):
$T_{\rm m}=5.6\times 10^{9}~{}[{\rm K}]\left(\frac{\Gamma_{\rm m}}{175}\right)^{-1}\left(\frac{\rho}{10^{14}~{}{\rm g~{}cm^{-3}}}\right)^{1/3}\left(\frac{Y_{\rm e}}{0.5}\right)^{1/3}\left(\frac{x_{\rm a}}{0.1}\right)^{1/3}\left(\frac{Z}{28}\right)^{5/3}.$ (A1)
We adopt $\Gamma_{\rm m}=175$ and $x_{\rm a}=0.1$ as fiducial values (Suwa, 2014). We consider nuclei with the magic neutron numbers $Z=28$, $50$, and $82$ (Haensel, 2001; Chamel & Haensel, 2008).
To estimate the time scale for crystallization, we continue the simulation in model BnMm2R10 until the temperature reaches the melting temperature defined by Equation (A1). The details of the setup are as follows. To ensure sufficient spatial resolution while maintaining reasonable computational cost, we conduct the simulation in three successive stages. We first set $r_{\rm out}$ to $20~{}{\rm km}$ with the number of numerical grid points of $N_{r}=16384$. Physical quantities at each grid point are initialized by linearly interpolating the data from model BnMm2 at time $t\approx 0.0354~{}{\rm s}$. For grid points located inside the innermost radial point of model BnMm2, we assign the values of the physical quantities at the innermost point of model BnMm2. We run a simulation by $t=0.0356~{}{\rm s}$, and then the resulting profile is used as the initial condition for the next stage. This procedure is repeated for $r_{\rm out}=10.1~{}{\rm km}$ and $10.01~{}{\rm km}$. This approach ensures $\Delta r<H/10$.
Figure 8: Left: Radial profiles of the gas density (top) and the gas temperature (bottom) in the vicinity of the PNS surface. Right: Plots for the gas temperature against the gas density (black line). Crystallizing temperature in the cases of $Z=28$ (solid), $Z=50$ (dotted), $Z=82$ (dashed) are represented by red lines (see Equation A1). All plots are measured at $t\approx 0.0357~{}{\rm s}$.
The left panels of Figure 8 show the radial profiles of the gas density (top) and temperature (bottom) in the vicinity of the PNS surface at $t\approx 0.0357~{}{\rm s}$. The gas density reaches $\sim 10^{15}~{}{\rm g~{}cm^{-3}}$ at the innermost grid, while the temperature remains nearly constant with radius. The neutrino cooling via the URCA process efficiently works at the innermost grid due to its high density.
The right panel of Figure 8 investigates the condition for the crystallization. The black line displays the relation between $\rho$ and $T$ at $t\approx 0.0357~{}{\rm s}$, while the red lines represent $T_{\rm m}$ for $Z=28$ (solid), $Z=50$ (dotted), and $Z=82$ (dashed). The figure indicates that the presented model satisfies the crystallization condition if $Z=82$. We define $t_{\rm crust}$ as the time at which the model satisfies the condition for $Z=82$ ($t_{\rm crust}\approx 0.0357~{}{\rm s}$).
The resulting time scale of the new crust formation is $t_{\rm crust}-t_{\rm stall}\sim 10^{-3}~{}{\rm s}$, which is roughly equal to $t_{\rm stall}^{\prime}$. Therefore, we conclude that the crystallization time scale does not significantly affect Figure 7 within the framework adopted in this study.